Mr. Kinders has contributed $78 at the end of each month into an RRSP paying 3...

Mr.

Kinders

has contributed

$78

at the end of each

month

into an RRSP paying

3 %

per annum compounded annually.

How much will Mr. Kinders have in the RRSP after 2020 year?

How much of the above amount is​ interest?
Answer #1

Monthly Contributions = $ 78, Contribution Tenure = 20 years or (20 x 12) = 240 months, Interest Rate = 3 %, Compounding Frequency: Monthly, Applicable Monthly Rate = 3/12 = 0.25 %

Total Future Value of Monthly Contributions (at the end of Year 20) = FV = 78 x (1.0025)^(239) + 78 x (1.0025)^(238) +............+ 78 = 78 x [{(1.0025)^(240)-1}/{(1.0025)-1}] = $ 25607.55585 ~ $ 25607.56

Interest Portion = FV - Number of Contributions x Value of Each Contribution = 25607.56 - 240 x 78 = $ 6887.5558 ~ $ 6887.56